# HierLoc: Hyperbolic Entity Embeddings for Hierarchical Visual Geolocation

### v1

@@ 1 Introduction
To encode the geographic hierarchy effectively, we represent entities in hyperbolic space. In Euclidean space, hierarchical structures become increasingly compressed as depth grows: the number of entities expands roughly exponentially from country → region → subregion → city, but Euclidean distances grow only linearly. This mismatch causes deeper-level entities to crowd together, reducing discriminative power during inference. Hyperbolic geometry, by contrast, naturally provides exponential volume growth and therefore allocates proportionally more space to represent large branching factors in deep hierarchies (Nickel and Kiela, 2017; Chen et al., 2021). As a result, related entities can remain close, while fine-grained locations can still be well separated, making hyperbolic space a more faithful and expressive embedding space for geographic hierarchies. This effect can be observed visually in the Figure 2 in the Appendix A.2. To our knowledge, this is the first application of Hyperbolic embeddings to represent hierarchical geographic entities for geolocation explicitly. In addition to expected continental clusters, Figure 6 in the Appendix shows the semantic relationships between entities directly leveraging strong semantics of modern image encoders such as DINOV3 (siméoni2025DINOV3). To couple this geometry with geographic structure, we introduce a Geo-Weighted Hyperbolic InfoNCE (GWH-InfoNCE) loss that weights negative logits using the haversine formula. This objective materially improves fine-scale discrimination while preserving global structure.
@@ 2 Related Works
Classical work framed geolocation as image retrieval against large galleries, e.g. IM2GPS (Hays and Efros, 2008), later revisited with stronger deep learning baselines (Vo et al., 2017). An alternative line casts the task as geocell classification, as in PlaNet (Weyand et al., 2016) and CPlaNet (Seo et al., 2018), while recent methods emphasize scalability (SC retrieval (Haas et al., 2023), PIGEON (Haas et al., 2024)) or generative modeling of geodesic uncertainty (Dufour et al., 2025). This trend continues with the work of LocDiff (Wang et al., 2025) with multi-scale latent diffusion. Benchmarks such as MediaEval’16 and OSV5M (Astruc et al., 2024) further standardized evaluation. The emergence of foundational models also resulted in works such as GeoReasonser (Li et al., 2025) and Img2Loc (Zhou et al., 2024), which push the boundary forward. Several hybrid approaches combine retrieval and classification, such as Translocator (Pramanick et al., 2022) and GeoDecoder (Clark et al., 2023), but these methods generally assume Euclidean label spaces. Although hierarchy is central to these protocols (e.g., country/region/city splits, quadtrees), prior work has used hierarchy mainly for evaluation or partitioning. By contrast, we treat the hierarchy itself as the object of representation, aligning images directly to entities across levels.
@@ 4.2 Ablations
The final ablation of removing the text and location modalities does reduce the performance a little across all levels, but it is not significant, and the image signal is the most important modality for this task. Together, these results validate our design choices: each component contributes meaningfully, and their combination yields the strongest performance. Furthermore, other ablations such as role of mean image embeddings, hyper parameter sensitivity, curvature choice for the Lorentz model and the choice of weight decaying function for the geo-weights in the loss are reported in the Appendix A.5. Further analysis of the computational efficiency of our framework against retrieval and egenrative methods is also reported in the Appendix A.8, which quantifies the improvements over existing retrieval methods.
@@ 5 Discussion
Several limitations remain. On datasets such as IM2GPS, YFCC4K, and IM2GPS3K, the absence of dense street-level entities constrains performance at the 1 km threshold. Our evaluation in Appendix A.9 further demonstrates difficulties in performance generalization between regions with less number of training samples. Looking ahead, the broader promise of HierLoc lies in its generality. Any task with hierarchically structured data such as taxonomies in biodiversity, linguistic families, or knowledge graphs could benefit from the same principles of Hyperbolic entity embeddings, multimodal fusion, and structured retrieval. Visual geolocation thus serves as a challenging and high-impact testbed, but the underlying methodology extends beyond it.
@@ 6 Ethics Statement
This work relies exclusively on publicly available datasets for visual geolocation, including OSV5M, MediaEval’16, YFCC4K, IM2GPS, and IM2GPS3K, which do not contain personally identifiable information beyond image content already released for research purposes. No additional human subjects were involved, and no private or sensitive data were collected.
We acknowledge that geolocation technologies may pose privacy and security risks if misapplied, for example in surveillance, tracking individuals, or identifying sensitive locations. Our work is intended solely for scientific benchmarking and methodological advancement in large-scale representation learning. All datasets are cited from their original sources and used under their research licenses.
We encourage future research to consider fairness and bias issues, particularly regarding underrepresented geographic regions, and to assess societal impacts of deploying such models.
@@ 7 Reproducibility Statement
We will release the source code in an anonymous repository for the camera-ready version of the paper, should it be accepted. This will include entity construction scripts, training pipelines, and evaluation scripts to enable exact replication of our results.
@@ A.2 Preliminaries on Hyperbolic Geometry
where $\|v\|$ is the Euclidean norm of $v$, and $\vec{x}\in\mathbb{R}^{d}$ are the "spacelike" coordinates of $x$. Note that $\log_{O}(x)$ is not simply the Euclidean projection of $\vec{x}$; the prefactor ensures exact agreement with the Riemannian logarithm. Figure 2 illustratively shows the process of projecting points from $\mathbb{R}^{d}$ to $\mathbb{H}_{K}^{d}$ using $\exp_{O}$ function, conversely, the process of projecting points from $\mathbb{H}_{K}^{d}$ to $\mathbb{R}^{d}$ using $\log_{0}$ function.

### v2

# arXiv:2601.23064v2[cs.CV] 02 Mar 2026

@@ 1 Introduction
To encode the geographic hierarchy effectively, we represent entities in hyperbolic space. In Euclidean space, hierarchical structures become increasingly compressed as depth grows: the number of entities expands roughly exponentially from country → region → subregion → city, but Euclidean distances grow only linearly. This mismatch causes deeper-level entities to crowd together, reducing discriminative power during inference. Hyperbolic geometry, by contrast, naturally provides exponential volume growth and therefore allocates proportionally more space to represent large branching factors in deep hierarchies (Nickel and Kiela, 2017; Chen et al., 2021). As a result, related entities can remain close, while fine-grained locations can still be well separated, making hyperbolic space a more faithful and expressive embedding space for geographic hierarchies. This effect can be observed visually in the Figure 2 in the Appendix A.2. To our knowledge, this is the first application of Hyperbolic embeddings to represent hierarchical geographic entities for geolocation explicitly. To couple HYperbolic geometry with geographic structure, we introduce a Geo-Weighted Hyperbolic InfoNCE (GWH-InfoNCE) loss that weights negative logits using the haversine formula. This objective materially improves fine-scale discrimination while preserving global structure.
@@ 2 Related Works
Classical work framed geolocation as image retrieval against large galleries, e.g. IM2GPS (Hays and Efros, 2008), later revisited with stronger deep learning baselines (Vo et al., 2017).Early pioneers like PlaNet (Weyand et al., 2016) and CPlaNet (Seo et al., 2018) used S2 cells to discretize the globe into multi-scale classes. Recent iterations like Where We Are and What We’re Looking at Clark et al. (2023) and PIGEON (Haas et al., 2024) have refined this via massive-scale classification and semantic fusion. While recent methods emphasize scalability (SC retrieval (Haas et al., 2023), PIGEON (Haas et al., 2024)) or generative modeling of geodesic uncertainty (Dufour et al., 2025). This trend continues with the work of LocDiff (Wang et al., 2025) with multi-scale latent diffusion. Benchmarks such as MediaEval’16 and OSV5M (Astruc et al., 2024) further standardized evaluation. The emergence of foundational models also resulted in works such as GeoReasonser (Li et al., 2025) and Img2Loc (Zhou et al., 2024), which push the boundary forward. Several hybrid approaches combine retrieval and classification, such as Translocator (Pramanick et al., 2022) and GeoDecoder (Clark et al., 2023).
While existing geo-classification methods use multi-scale cells for partitioning, they often collapse into a single-level output, losing critical hierarchical signals. Our method explicitly preserves and leverages this hierarchical structure during both representation and inference.
@@ 4.2 Ablations
The ablation of removing the text and location modalities shown in the table 7 does reduce the performance a little across all levels, but it is not significant, and the image signal is the most important modality for this task. But the combination of all three modalities results in the best performance. Furthermore, other ablations such as role of mean image embeddings, hyper parameter sensitivity, curvature choice for the Lorentz model and the choice of weight decaying function for the geo-weights in the loss are reported in the Appendix A.5. Further analysis of the computational efficiency of our framework against retrieval and generative methods is also reported in the Appendix A.8, which quantifies the improvements over existing retrieval methods. We perform an ablation study to evaluate the impact of the hierarchical structure by comparing our approach against a flat retrieval baseline (Table 8). We further analyze the effect of beam search width on localization performance. Figure 5 illustrates the beam search inference process on a geographic map, demonstrating how candidate entities are identified and refined across different hierarchical levels using a beam width of 10.
@@ 5 Discussion
Several limitations remain, 1 km localization performance of our method is inherently bounded by a discretization limit resulting from our selection of 240k fixed entities, which lacks dense street-level granularity. We acknowledge this as a fundamental trade-off: our hierarchical representation prioritizes structural and regional interpretability over sub-kilometer accuracy. Our evaluation in Appendix A.9 further demonstrates difficulties in performance generalization between regions with less number of training samples.
Looking ahead, the broader promise of HierLoc lies in its generality. Any task with hierarchically structured data such as taxonomies in biodiversity, linguistic families, or knowledge graphs could benefit from the same principles of Hyperbolic entity embeddings, multimodal fusion, and structured retrieval. Visual geolocation thus serves as a challenging and high-impact testbed, but the underlying methodology extends beyond it.
@@ 6 Ethics Statement
This work relies exclusively on publicly available datasets for visual geolocation, including OSV5M, MediaEval’16, YFCC4K, IM2GPS, and IM2GPS3K, which do not contain personally identifiable information beyond image content already released for research purposes. No additional human subjects were involved, and no private or sensitive data were collected. We acknowledge that geolocation technologies may pose privacy and security risks if misapplied, for example in surveillance, tracking individuals, or identifying sensitive locations. Our work is intended solely for scientific benchmarking and methodological advancement in large-scale representation learning. All datasets are cited from their original sources and used under their research licenses. We encourage future research to consider fairness and bias issues, particularly regarding underrepresented geographic regions, and to assess societal impacts of deploying such models.
@@ A.2 Preliminaries on Hyperbolic Geometry
where $\|v\|$ is the Euclidean norm of $v$, and $\vec{x}\in\mathbb{R}^{d}$ are the ”spacelike” coordinates of $x$. Note that $\log_{O}(x)$ is not simply the Euclidean projection of $\vec{x}$; the prefactor ensures exact agreement with the Riemannian logarithm. Figure 2 illustratively shows the process of projecting points from $\mathbb{R}^{d}$ to $\mathbb{H}_{K}^{d}$ using $\exp_{O}$ function, conversely, the process of projecting points from $\mathbb{H}_{K}^{d}$ to $\mathbb{R}^{d}$ using $\log_{0}$ function.
@@ A.9 Qualitative Analysis
Figure 5: Hierarchical beam search visualization for visual geolocation. Given a query street-view image (a), HierLoc performs a coarse-to-fine geographic search through four hierarchical levels (b). At each level, the model scores candidate regions and selects the highest-scoring entity (red path). Panels (c)–(e) show the candidate regions at each level as a choropleth, where color intensity indicates the model’s confidence score normalized within that level. Panel (f) shows the final city-level prediction on an OpenStreetMap basemap, with the predicted location (red cross) and the ground truth image coordinate (gold star).
To provide interpretable insight into the inference behavior of HierLoc, we visualize the full hierarchical beam search trace alongside its geographic realization across the four entity levels in the figure 5. Panel (b) displays the beam search tree, where each column corresponds to a hierarchy level (Country, Region, Sub-region, City) and each node represents a candidate geographic entity scored by computing the squared hyperbolic distance between the query image embedding and the entity embedding in the Lorentz manifold. The predicted path, shown in red, traces the highest-scoring entity selected at each level. Pruned candidates appear as faint background nodes, illustrating the breadth of the search space that the beam width of k=10 efficiently narrows at each step. Panels (c) through (f) project this search onto geographic maps with progressively increasing zoom, mirroring the coarse-to-fine structure of the hierarchy. At the country level (c), all 10 candidate countries are rendered on a world map with choropleth shading proportional to their normalized scores. At the region level (d), the view restricts to regions within the selected country, and similarly for sub-regions (e) within the selected region. This geographic drill-down directly reflects the computational savings of hierarchical beam search, which reduces the candidate space from 240k entities to a tractable subset at each level. The city-level panel (f) overlays the final predicted location and ground truth coordinate on an OpenStreetMap basemap, with a dotted line and distance annotation quantifying the localization error. In this example, HierLoc correctly identifies Peru, narrows to the Cusco region, localizes to Provincia de Urubamba, and predicts Ollantaytambo within 11.0 km of the true coordinate, demonstrating the model’s ability to perform structured geographic reasoning through the learned hyperbolic entity hierarchy.